OpenStudy (anonymous):

find all solutions of the equation tan3x+1=sec3x in the interval [0,2pi).

OpenStudy (amistre64):

is that \(\tan^3x\) or \(\tan3x\)?

OpenStudy (anonymous):

3x

OpenStudy (amistre64):

so if we clean this up we have: tan u + 1 = sec u

OpenStudy (amistre64):

if we square both sides that might help us, but it tends to introduce a few false results to be weeded out

OpenStudy (amistre64):

tan^2 + 1 + 2tan = sec^2 , but sec^2 = tan^2 + 1 tan^2 + 1 + 2tan = tan^2 + 1 2tan = 0 might be worth looking into

OpenStudy (amistre64):

well, thats one of them :)

OpenStudy (amistre64):

hmm, might be the only one too